# Multi-Constraint Safe RL with Objective Suppression for Safety-Critical Applications

### v2

@@ 5.2 Results
We compare our method with Recovery RL. Each method is evaluated using 50 rollouts from the policy in different environments. The results are shown in Table 2. Compared with the baseline Recoevery RL, our method outperforms on constraint satisfaction while only incurring a small decrease in task reward.
@@ 6 Conclusion
We propose Objective Suppression, a novel algorithm that adaptively suppresses the task reward objective to enforce safety constraints. Combined with existing safe RL algorithms, we demonstrate that Objective Suppression can maintain the task reward of the base algorithm while significantly lowering the constraint violations in multi-constraint scenarios, including an autonomous driving domain where incorrect behaviors can be disastrous.

### v3

@@ 5.2 Results
We compare our method with Reward Penalty and Recovery RL. Each method is evaluated using 50 rollouts from the policy in different environments. The results are shown in Table 2. Compared with the baseline Recovery RL, our method outperforms on constraint satisfaction while only incurring a small decrease in task reward. Compared with Recovery RL, our method reduces collision to 32.5% and out-of-lane violations to 39.1%.
Proprietary Data and Real World
We also demonstrate a variant of this approach on extensively collected proprietary driving datasets containing thousands of real-world miles covering a variety of challenging conditions. On these datsets, we observe an increase in distance traveled (a proxy for success) combined with a $2\%$ improvement in collision rate, and a $38\%$ improvement in harsh braking events compared to Recovery RL alone. This evaluation in simulation enabled us to deploy our method on a fleet of vehicles in the real world with performance mirroring the results in simulation.
@@ 6 Conclusion
We identified that the widely-used CMDP model for safe RL can potentially allow dangerous behaviors in less frequently visited states. We proposed UCMDPs that uniformly constrain reachable states to counter this issue. To solve UCMDPs, we propose Objective Suppression, a novel algorithm that adaptively suppresses the task reward objective to enforce safety constraints. Combined with existing safe RL algorithms, we demonstrate that Objective Suppression can maintain the task reward of the base algorithm while significantly lowering the constraint violations in multi-constraint scenarios, including an autonomous driving domain where incorrect behaviors can be disastrous.
7 Limitations
In our evaluations, we observe that $\kappa$ is still an important tunable hyperparameter for Objective Suppression. It would be interesting to conduct a more thorough study on how this parameter impacts the task reward-risk trade-off. On the other hand, both our benchmarks are continuous control domains, despite our method also applies to discrete action spaces. We believe evaluating Objective Suppression in more domains with discrete action spaces remains an interesting future direction.